Let X1, . . . , Xn ~(iid) Bernoulli(p), and let Pn=n-1(Xn+ ... + Xn) .

(a) Give an exact expression for  P.5<Pn).

b) Evaluate your expression from part (a) for n = 200 and p = 4/9.
Answer #1

X1, 72, ..., Xu be iid Bernoulli cp). Pn = X1 + Xq t... + xw w (a) To find p(0.5<bn) = P(Pm 30.5) = p/ Xgt -. . t Xnxos) = Pl